# Edinburgh 2025/29

### v1

@@ Abstract
We present a global determination of parton distribution functions (PDFs) that accounts for higher twist corrections in deep-inelastic scattering (DIS) and linear power corrections for single inclusive jet and dijet production data from the LHC. We determine these corrections and their associated correlated uncertainties using a methodology based on the theory covariance formalism, previously used to account for nuclear uncertainties and missing higher order uncertainties (MHOUs) in global PDF determinations. We then study the impact of the power corrections on the extracted PDFs, and demonstrate an improved description of the datam due to a reduced sensitivity to DIS data in the low-$x$ region where higher twist uncertainties are relatively large, and a reduced sensitivity to single inclusive jet data at relatively low $p_{T}$, where linear power corrections can be significant. Finally, we assess the impact of power corrections on observables relevant to LHC phenomenology, including Higgs production via gluon fusion, and the determination of $\alpha_{s}$. We find that these effects, while small, can be significant, improving perturbative convergence.
@@ 1 Introduction
The advent of the high energy frontier, now at a center-of-mass energy scale of 13.6 TeV at the LHC, offers an unprecedented opportunity to test the predictive power of the Standard Model. Remarkable progress has been made in calculations at next-to-next-to-leading order (NNLO) and even next-to-next-to-next-to-leading order (N3LO) accuracy in perturbative QCD, together with the inclusion of mass corrections, QED and Electroweak corrections. Nevertheless, some significant sources of theoretical uncertainty remain, both perturbative and non-perturbative, and in both the hard cross-sections themselves and in the parton distribution functions (PDFs). If our theoretical predictions are to match the accuracy and precision of LHC data, we need to control all theoretical uncertainties at the one percent level [Amoroso:2022eow].
@@
Theoretical uncertainties arising from missing higher-orders (MHOUs) have been included in PDF fits using a theory covariance matrix [Ball:2018lag] constructed from scale variations [NNPDF:2019vjt, NNPDF:2019ubu, NNPDF:2024dpb], or by fitting nuisance parameters [McGowan:2022nag]. The covariance matrix approach has also been used to incorporate the non-perturbative uncertainties due to nuclear effects in global fits with deep-inelastic scattering (DIS) and Drell-Yan (DY) data on heavy nuclei and deuteron targets [Ball:2018twp, Ball:2020xqw].
@@
Besides theoretical uncertainties due to perturbative corrections and nuclear effects, there are further uncertainties due to effects which are suppressed by a power of the hard scale. In DIS, the familiar expression for the cross-section, factorised into a convolution of a perturbative coefficient function and universal PDFs, is only the leading term in the operator-product expansion. Additional terms, known as higher twist (HT) corrections, are always present, but suppressed by powers of the hard scale $Q^{2}$. These HT corrections are in essence non-perturbative, and over the years there have been many attempts to determine them empirically by fitting $\mathcal{O}(\Lambda^{2}/Q^{2})$ corrections (here $\Lambda$ is the QCD scale) in the context of a PDF fit [Virchaux:1991jc, Alekhin:1998df, Alekhin:1999af, Martin:1998np, Martin:2003sk, Blumlein:2008kz, Alekhin:2012ig, Ball:2013gsa, Thorne:2014toa, Alekhin:2017kpj, Alekhin:2022tip, Cerutti:2025yji]. This is challenging, as the HT corrections can easily be confused with the unknown terms behaving as powers of $\mathcal{O}\left(1/\log(Q^{2}/\Lambda^{2})\right)$ arising due to MHOUs. Consequently most global PDF fits impose kinematic cuts on DIS data to ensure that the HT corrections are small.
@@
For hadronic processes, such as DY or inclusive jets, there is no operator product expansion, but factorization still holds in the limit of large momentum transfer [Collins:1989gx]. Again there are non-perturbative contributions suppressed by powers of the hard scale $Q_{T}$. For sufficiently inclusive processes with a color singlet final state, such as DY, the power corrections (PCs) are expected to be $\mathcal{O}(\Lambda^{2}/Q_{T}^{2})$ [Beneke:1995pq, Caola:2021kzt], and thus small except close to threshold. However for many other hadronic observables there are generally linear power corrections of $\mathcal{O}(\Lambda/Q_{T})$. For example, linear PCs to jet processes have been studied extensively in $e^{+}e^{-}$ annihilation [Dasgupta:2003iq, Dasgupta:2007wa, Dasgupta:2008di, Cacciari:2007aq], with analytic estimates of the shift in the jet energy due to hadronization and underlying event. Studies of jets in hadronic collisions [Olness:2009qd] have extended these analyses to estimate the size of theoretical uncertainties related to various non-perturbative effects in inclusive single jet cross-sections. These estimates suggest that the linear PCs to jet processes might be significant even up to scales as large as $100$ GeV. Linear PCs are also expected in top production [Makarov:2023uet], unless one uses a short distance scheme for the top quark mass. Since linear PCs fall off much more slowly with hard scale than HT, it is even more difficult to disentangle them from MHOUs. To the best of our knowledge, an empirical determination of linear PCs to jets in the context of a global PDF fit has not yet been attempted.
@@
In this paper, we use the theory covariance matrix formalism developed in [Ball:2018lag, Ball:2021icz], recently validated in Ref. [Ball:2025xgq], to provide reliable estimates of the corrections and theoretical uncertainties due to HT in DIS, and to PCs in single jet inclusive cross-sections and dijet cross-sections. The analysis is performed entirely within the open-source NNDPF framework [NNPDF:2021uiq]. The paper is organised as follows. In section 2 the methodology is discussed in detail. In section 3 we will present results, firstly for the HT, and then for PCs to inclusive jets and dijets. In section 4, we will include these corrections and uncertainties in a global PDF fit, assess its impact on the PDFs and relative luminosities, on the theoretical predictions for fundamental LHC processes such as the Higgs boson production, and on the determination of the strong coupling $\alpha_{s}$ from hadronic data. Finally, in section 5 we summarise our main conclusions and provide information on how to access the PDF fits with HT and PCs in LHAPDF format [Buckley:2014ana].
@@ 2 Methodology for the Determination of Power Corrections
In this section, we review the theory covariance methodology we will use to estimate the theoretical uncertainties associated with power corrections. This is based on the general formalism developed in Ref. [Ball:2021icz], where it is shown how to account for correlations between theory uncertainties in the PDFs and in the matrix element used in a prediction. When applied to parameters of physical interest, the methodology allows for the extraction of such parameters, together with their correlated uncertainties, in such a way as to account for all the correlations with the underlying PDFs. As such, the methodology was recently validated in a determination of the strong coupling [Ball:2025xgq], by comparing results to those obtained using the more cumbersome correlated replica method [Ball:2018iqk], and by closure testing [NNPDF:2017mvq]. In the following, we provide a brief overview of the theory covariance methodology before moving to its specific application to power corrections.
@@ 2.1 The Theory Covariance Method
The accurate and precise determination of PDFs requires careful consideration of the theoretical uncertainties associated with theoretical predictions. Theoretical uncertainties and their correlations can be incorporated in the fitting procedure through the implementation of a theoretical covariance matrix $C^{\textrm{th}}$. If we assume that these uncertainties are Gaussian and independent of the experimental uncertainties, the total covariance matrix $C$ can be written as the sum of the experimental $t_{0}$ covariance matrix [Ball:2009qv] $C^{\textrm{exp}}_{t_{0}}$ and a theory covariance matrix [Ball:2018lag, Ball:2018twp, Ball:2020xqw, NNPDF:2019vjt, NNPDF:2019ubu, NNPDF:2024dpb] incorporating in particular all MHOUs and nuclear uncertainties
@@
where $T_{i}$ is the theory prediction for data point $i$. The figure of merit in Eq. (2.3) ensures that the relative weights of the data points during the fit account for both experimental and theoretical uncertainties [Ball:2018lag]. The $\chi^{2}$ in Eq. (2.3) is then optimised for each replica $k$ in the Monte Carlo ensemble, using a cross-validation procedure to prevent overfitting. In terms of Bayesian reasoning, minimization of the $\chi^{2}$ corresponds to the maximization of the conditional probability of the predictions given the data
@@
Following Ref. [Ball:2021icz], additional sources of theoretical uncertainty can be modelled as fully correlated shifts in the predictions induced by a set of $N_{\textrm{nuis}}$ sources of theoretical uncertainty
@@
The theory covariance matrix $S$ to be used in the fit will then be constructed as in Eq. (2.7) using the shifts in Eq. (2.16). Note that since the corrections in Eq. (2.15) are linear in the nuisance parameters, there is no need (unlike in Refs. [AbdulKhalek:2019bux, AbdulKhalek:2019ihb, NNPDF:2024dpb, Ball:2025xgq]) to consider contributions from positive and negative shifts separately, nor to consider complicated prescriptions: Eq. (2.7) is perfectly sufficient. After performing the fit including the new theory covariance matrix, we can compute the posterior for the nuisance parameters using Eq. (2.11), and their covariance using Eq. (2.12), which after the rescaling (2.16) gives the posterior parameters $h^{\rm post}_{\alpha}$ and their covariance
@@ 2.2 Modelling Higher Twist Corrections in DIS
There are two distinct approaches to modelling higher twist corrections to theoretical predictions for DIS data: the additive approach [Alekhin:1998df, Alekhin:1999af, Blumlein:2008kz, Alekhin:2012ig, Alekhin:2017kpj, Alekhin:2022tip], and the multiplicative approach [Virchaux:1991jc, Martin:1998np, Martin:2003sk, Ball:2013gsa, Thorne:2014toa]. In the additive approach one ignores all perturbative corrections to the HT. In the multiplicative approach, one first factors out the leading twist contribution, so in effect one assumes that the HT evolves perturbatively in the same way as the leading twist. Both approaches are thus approximate, and which one adopts is largely a matter of taste (see Ref. [Cerutti:2025yji] for a recent phenomenological comparison).
We consider HT corrections as multiplicative shifts to DIS observables, which comprise measurements of structure functions $F_{2}$ and reduced cross-sections. We model the shifts as functions of Bjorken-$x$, to be determined empirically. Moreover, given that in principle HT can depend on nuclear effects, the shifts for proton and deuteron targets are parametrised independently [Cerutti:2025yji]. Hence, we write the measured structure functions as
@@
For ratio data, we have
@@
Finally, DIS CC cross-sections are corrected similarly, by introducing a common multiplicative shift for CHORUS, NuTeV, and HERA CC data regardless of the nuclear target used in the experiment. Thus we will assume that nuclear effects at higher twist in CC data are similar to those included at leading twist [Ball:2018twp]. This approximation will be justified a priori. Hence, the shift for any CC cross-section will be written as
@@ 2.3 Modelling Linear Power Corrections for Jet and Dijet Cross-sections
Here $f_{p}(x_{1})$ and $f_{q}(x_{2})$ are the PDFs of the incoming partons $p$ and $q$, and $d\hat{\sigma}_{pq\rightarrow\textrm{jet}}/dp_{T}dy$ is the partonic cross-section computed at NNLO in perturbative QCD. MHOUs estimated from scale variations have been accounted for in a variant of NNPDF4.0 fits [NNPDF:2024dpb]. While it has been understood for some time that there are linear power corrections to these processes, from a wide variety of sources, in global PDF fits these are generally ignored.
@@
There are many factors that influence the measurement of jet cross-sections, mainly to do with the way the jets and their associated energy scales are defined. For instance, the algorithm used to identify the jets introduces a dependence on the radius of the jet $R$. Furthermore, assuming that jets have been faithfully reconstructed, there is a variety of physical phenomena, not all belonging to the domain of perturbative calculations, that can affect the measured energy of the jet. In other words, the energy of the jet does not necessarily match the energy of the hard parton that initiated the jet. For example, the underlying event originates from uncontrolled interactions in the initial proton-proton collision (multiple parton interactions and colour interactions), as well as contributions from initial- and final-state radiation. As a results, the measured energy of the jet is shifted due to the external contribution from the underlying event that flows into the jet. Investigations to disentangle the underlying event contribution from single-inclusive jets exploiting Monte Carlo event generators have been performed by ATLAS [ATLAS:2014iez] and CMS [CMS:2013ycn]. On the other hand, in the process of hadronization, some of the jet energy leaks out from the jet cone, reducing the energy that is eventually measured. This process is modelled through the use of final state Monte Carlo generators. Experimentalists correct for these effects in their measurements, with associated systematic uncertainties.
@@
However in addition to all this, the theoretical calculation of the perturbative jet cross-section is also subject to non-perturbative uncertainties, resulting in power corrections of $O(\Lambda/p_{T})$ to the parton level transverse momentum and thus to the perturbative cross-sections. These genuine non-perturbative effects are by their nature very difficult to disentangle from the quasi-partonic effects already corrected for in the experimental measurements and thus, as with higher twist, an empirical approach is the only practical way to assess them.
@@
Whatever the source of the effect, a change in energy results in a shift $\delta p_{T}$ to the fixed parton level transverse momentum $p_{T}$ of the jet. The correction $\delta p_{T}$ has been estimated in Refs. [Dasgupta:2007wa, Dasgupta:2008di, Cacciari:2007aq]. Since jet cross-sections fall very steeply with $p_{T}$ (typically as $\sim p_{T}^{-n}$, with $n$ as large as six or seven), such a shift results in a multiplicative power correction of the form $1+n\delta p_{T}/p_{T}$, which may be large simply because $n$ is large [Olness:2009qd].
@@
We will thus model the shift in the single-inclusive jet cross-section induced by power corrections as
@@
Note that since the linear power corrections fall off much more slowly with the hard scale than the HT corrections to DIS, they will be even more difficult to disentangle from MHO perturbative corrections. This problem is compounded by the fact that the N3LO corrections to jet processes are unknown, to the perturbative predictions are less precise than for DIS. On the other hand, since as explained in [Olness:2009qd] the scale of the linear power corrections may be relatively large (several GeV rather than several hundred MeV), they may continue to be significant up to relatively high scales. They can thus be significant, even at the LHC.

### v2

@@ Abstract
We present a global determination of parton distribution functions (PDFs) that accounts for higher twist corrections in deep-inelastic scattering (DIS) and linear power corrections for single inclusive jet and dijet production data from the LHC. We determine these corrections and their associated correlated uncertainties using a methodology based on the theory covariance formalism, previously used to account for nuclear uncertainties and missing higher order uncertainties (MHOUs) in global PDF determinations. We then study the impact of the power corrections on the extracted PDFs, and demonstrate an improved description of the data due to a reduced sensitivity to DIS data in the low-$x$ region where higher twist uncertainties are relatively large, and a reduced sensitivity to single inclusive jet data at relatively low $p_{T}$, where hadronization corrections can be large. Finally, we assess the impact of power corrections on observables relevant to LHC phenomenology, including Higgs production via gluon fusion, and the determination of $\alpha_{s}$. We find that these effects, while small, can be significant, improving perturbative convergence.
@@ 1 Introduction
The advent of the high energy frontier, now at a center-of-mass energy scale of 13.6 TeV at the LHC, offers an unprecedented opportunity to test the predictive power of the Standard Model. Remarkable progress has been made in calculations at next-to-next-to-leading order (NNLO) and even next-to-next-to-next-to-leading order (N3LO) accuracy in perturbative QCD, together with the inclusion of mass corrections, QED and Electroweak corrections. Nevertheless, some significant sources of theoretical uncertainty remain, both perturbative and non-perturbative, and in both the hard cross-sections themselves and in the parton distribution functions (PDFs). If our theoretical predictions are to match the accuracy and precision of LHC data, we need to control all theoretical uncertainties at the one percent level [16].
@@
Theoretical uncertainties arising from missing higher-orders (MHOUs) have been included in PDF fits using a theory covariance matrix [29] constructed from scale variations [5, 7, 35], or by fitting nuisance parameters [69]. The covariance matrix approach has also been used to incorporate the non-perturbative uncertainties due to nuclear effects in global fits with deep-inelastic scattering (DIS) and Drell-Yan (DY) data on heavy nuclei and deuteron targets [30, 31].
@@
Besides theoretical uncertainties due to perturbative corrections and nuclear effects, there are further uncertainties due to effects which are suppressed by a power of the hard scale. In DIS, the familiar expression for the cross-section, factorised into a convolution of a perturbative coefficient function and universal PDFs, is only the leading term in the operator-product expansion. Additional terms, known as higher twist (HT) corrections, are always present, but suppressed by powers of the hard scale $Q^{2}$. These HT corrections are in essence non-perturbative, and over the years there have been many attempts to determine them empirically by fitting $\mathcal{O}(\Lambda^{2}/Q^{2})$ corrections (here $\Lambda$ is the QCD scale) in the context of a PDF fit [75, 15, 13, 67, 66, 43, 11, 24, 73, 10, 14, 51, 63, 12]. This is challenging, as the HT corrections can easily be confused with the unknown terms behaving as powers of $\mathcal{O}\left(1/\log(Q^{2}/\Lambda^{2})\right)$ arising due to MHOUs. Consequently most global PDF fits impose kinematic cuts on DIS data to ensure that the HT corrections are small.
@@
For hadronic processes, such as DY or inclusive jets, there is no operator product expansion, but factorization still holds in the limit of large momentum transfer [54]. Again there are non-perturbative contributions suppressed by powers of the hard scale $Q_{T}$. For sufficiently inclusive processes with a color singlet final state, such as DY, the power corrections (PCs) are expected to be $\mathcal{O}(\Lambda^{2}/Q_{T}^{2})$ [41, 49], and thus small except close to threshold. However for many other hadronic observables there are generally linear power corrections of $\mathcal{O}(\Lambda/Q_{T})$. For example, linear PCs to jet processes have been studied extensively in $e^{+}e^{-}$ annihilation [58, 56, 57, 46], with analytic estimates of the shift in the jet energy due to hadronization and underlying event. Studies of jets in hadronic collisions [70] have extended these analyses to estimate the size of theoretical uncertainties related to various non-perturbative effects in inclusive single jet cross-sections. These estimates suggest that the linear PCs to jet processes might be significant even up to scales as large as $100$ GeV. Linear PCs are also expected in top production [65], unless one uses a short distance scheme for the top quark mass. Since linear PCs fall off much more slowly with hard scale than HT, it is even more difficult to disentangle them from MHOUs. To the best of our knowledge, an empirical determination of linear PCs to jets in the context of a global PDF fit has not yet been attempted.
@@
In this paper, we use the theory covariance matrix formalism developed in [29, 38], recently validated in Ref. [20], to provide reliable estimates of the corrections and theoretical uncertainties due to HT in DIS, and to PCs in single jet inclusive cross-sections and dijet cross-sections. The analysis is performed entirely within the open-source NNDPF framework [33]. The paper is organised as follows. In section 2 the methodology is discussed in detail. In section 3 we will present results, firstly for the HT, and then for PCs to inclusive jets and dijets. In section 4, we will include these corrections and uncertainties in a global PDF fit, assess its impact on the PDFs and relative luminosities, on the theoretical predictions for fundamental LHC processes such as the Higgs boson production, and on the determination of the strong coupling $\alpha_{s}$ from hadronic data. Finally, in section 5 we summarise our main conclusions and provide information on how to access the PDF fits with HT and PCs in LHAPDF format [45].
@@ 2 Methodology for the Determination of Power Corrections
In this section, we review the theory covariance methodology we will use to estimate the theoretical uncertainties associated with power corrections. This is based on the general formalism developed in Ref. [38], where it is shown how to account for correlations between theory uncertainties in the PDFs and in the matrix element used in a prediction. When applied to parameters of physical interest, the methodology allows for the extraction of such parameters, together with their correlated uncertainties, in such a way as to account for all the correlations with the underlying PDFs. As such, the methodology was recently validated in a determination of the strong coupling [20], by comparing results to those obtained using the more cumbersome correlated replica method [26], and by closure testing [32]. In the following, we provide a brief overview of the theory covariance methodology before moving to its specific application to power corrections.
@@ 2.1 The Theory Covariance Method
The accurate and precise determination of PDFs requires careful consideration of the theoretical uncertainties associated with theoretical predictions. Theoretical uncertainties and their correlations can be incorporated in the fitting procedure through the implementation of a theoretical covariance matrix $C^{\textrm{th}}$. If we assume that these uncertainties are Gaussian and independent of the experimental uncertainties, the total covariance matrix $C$ can be written as the sum of the experimental $t_{0}$ covariance matrix [28] $C^{\textrm{exp}}_{t_{0}}$ and a theory covariance matrix [29, 30, 31, 5, 7, 35] incorporating in particular all MHOUs and nuclear uncertainties
@@
where $T_{i}$ is the theory prediction for data point $i$. The figure of merit in Eq. (2.3) ensures that the relative weights of the data points during the fit account for both experimental and theoretical uncertainties [29]. The $\chi^{2}$ in Eq. (2.3) is then optimised for each replica $k$ in the Monte Carlo ensemble, using a cross-validation procedure to prevent overfitting. In terms of Bayesian reasoning, minimization of the $\chi^{2}$ corresponds to the maximization of the conditional probability of the predictions given the data
@@
Following Ref. [38], additional sources of theoretical uncertainty can be modelled as fully correlated shifts in the predictions induced by a set of $N_{\textrm{nuis}}$ sources of theoretical uncertainty
@@
The theory covariance matrix $S$ to be used in the fit will then be constructed as in Eq. (2.7) using the shifts in Eq. (2.16). Note that since the corrections in Eq. (2.15) are linear in the nuisance parameters, there is no need (unlike in Refs. [4, 6, 35, 20]) to consider contributions from positive and negative shifts separately, nor to consider complicated prescriptions: Eq. (2.7) is perfectly sufficient. After performing the fit including the new theory covariance matrix, we can compute the posterior for the nuisance parameters using Eq. (2.11), and their covariance using Eq. (2.12), which after the rescaling (2.16) gives the posterior parameters $h^{\rm post}_{\alpha}$ and their covariance
@@ 2.2 Modelling Higher Twist Corrections in DIS
There are two distinct approaches to modelling higher twist corrections to theoretical predictions for DIS data: the additive approach [15, 13, 43, 11, 10, 14], and the multiplicative approach [75, 67, 66, 24, 73]. In the additive approach one ignores all perturbative corrections to the HT. In the multiplicative approach, one first factors out the leading twist contribution, so in effect one assumes that the HT evolves perturbatively in the same way as the leading twist. Both approaches are thus approximate, and which one adopts is largely a matter of taste (see Ref. [51] for a recent phenomenological comparison).
We consider HT corrections as multiplicative shifts to DIS observables, which comprise measurements of structure functions $F_{2}$ and reduced cross-sections. We model the shifts as functions of Bjorken-$x$, to be determined empirically. Moreover, given that in principle HT can depend on nuclear effects, the shifts for proton and deuteron targets are parametrised independently [51]. Hence, we write the measured structure functions as
@@
Here we make the simplifying assumption that higher twist corrections to $F_{L}$ is the same as for $F_{2}$, since the data are not good enough to determine them independently. For ratio data, we have
@@
Finally, DIS CC cross-sections are corrected similarly, by introducing a common multiplicative shift for CHORUS, NuTeV, and HERA CC data regardless of the nuclear target used in the experiment. Thus we will assume that CC $F_{2}$, $F_{3}$ and $F_{L}$ are all corrected by the same factor, and furthermore that nuclear effects at higher twist in CC data are similar to those included at leading twist [30]. Hence, the shift for any CC cross-section will be written as
@@ 2.3 Modelling Linear Power Corrections for Jet and Dijet Cross-sections
Here $f_{p}(x_{1})$ and $f_{q}(x_{2})$ are the PDFs of the incoming partons $p$ and $q$, and $d\hat{\sigma}_{pq\rightarrow\textrm{jet}}/dp_{T}dy$ is the partonic cross-section computed at NNLO in perturbative QCD. MHOUs estimated from scale variations have been accounted for in a variant of NNPDF4.0 fits [35]. While it has been understood for some time that there are linear power corrections to these processes, from a wide variety of sources, in global PDF fits these are generally ignored.
@@
There are many factors that influence the measurement of jet cross-sections, mainly to do with the way the jets and their associated energy scales are defined. For instance, the algorithm used to identify the jets introduces a dependence on the radius of the jet $R$. Furthermore, assuming that jets have been faithfully reconstructed, there is a variety of physical phenomena, not all belonging to the domain of perturbative calculations, that can affect the measured energy of the jet. In other words, the energy of the jet does not necessarily match the energy of the hard parton that initiated the jet. For example, the underlying event originates from uncontrolled interactions in the initial proton-proton collision (multiple parton interactions and colour interactions), as well as contributions from initial- and final-state radiation. As a results, the measured energy of the jet is shifted due to the external contribution from the underlying event that flows into the jet. Investigations to disentangle the underlying event contribution from single-inclusive jets exploiting Monte Carlo event generators have been performed by ATLAS [3] and CMS [52]. On the other hand, in the process of hadronization, some of the jet energy leaks out from the jet cone, reducing the energy that is eventually measured. This process is modelled through the use of final state Monte Carlo generators. Experimentalists can correct for these effects in their measurements, by computing a correction factor with associated systematic uncertainties.
@@
However in addition to all this, the theoretical calculation of the perturbative jet cross-section is also subject to non-perturbative uncertainties, resulting in power corrections of $O(\Lambda/p_{T})$ to the parton level transverse momentum and thus to the perturbative cross-sections. These genuine non-perturbative effects are by their nature very difficult to disentangle from the quasi-partonic effects estimated by using final state Monte Carlos, and thus, as with higher twist, an empirical approach is the most practical way to assess them quantitatively.
@@
Whatever the source of the effect, a change in energy results in a shift $\delta p_{T}$ to the fixed parton level transverse momentum $p_{T}$ of the jet. The correction $\delta p_{T}$ has been estimated in Refs. [56, 57, 46]. Since jet cross-sections fall very steeply with $p_{T}$ (typically as $\sim p_{T}^{-n}$, with $n$ as large as six or seven), such a shift results in a multiplicative power correction of the form $1+n\delta p_{T}/p_{T}$, which may be large simply because $n$ is large [70].
@@
We will thus model the shift in the single-inclusive jet cross-section induced by all these various power corrections as
@@
Note that since the linear power corrections fall off much more slowly with the hard scale than the HT corrections to DIS, they will be even more difficult to disentangle from MHO perturbative corrections. This problem is compounded by the fact that the N3LO corrections to jet processes are unknown, to the perturbative predictions are less precise than for DIS. On the other hand, since as explained in [70] the scale of the linear power corrections may be relatively large (several GeV rather than several hundred MeV), they may continue to be significant up to relatively high scales. They can thus be significant, even at the LHC.